# Exploring the ultra-hot Jupiter WASP-178b

### v2

@@
Despite recent progress in the spectroscopic characterization of individual exoplanets, the atmospheres of key ultra-hot Jupiters (UHJs) still lack comprehensive investigations. These include WASP-178b, one of the most irradiated UHJs known to date. We observed the dayside emission signal of this planet with CRIRES+ in the spectral K-band. By applying the cross-correlation technique and a Bayesian retrieval framework to the high-resolution spectra, we identified the emission signature of 12CO (S/N = 8.9) and H2O (S/N = 4.9), and a strong atmospheric thermal inversion. A joint retrieval with space-based secondary eclipse measurements from TESS and CHEOPS allows us to refine our results on the thermal profile and thus to constrain the atmospheric chemistry, yielding a solar to super-solar metallicity ( dex) and a solar C/O ratio (). We infer a significant excess of spectral line broadening and identify a slight Doppler-shift between the 12CO and H2O signals. These findings provide strong evidence for a super-rotating atmospheric flow pattern and suggest the possible existence of chemical inhomogeneities across the planetary dayside hemisphere. In addition, the inclusion of photometric data in our retrieval allows us to account for stellar light reflected by the planetary atmosphere, resulting in an upper limit on the geometric albedo (0.23). The successful characterization of WASP-178b’s atmosphere through a joint analysis of CRIRES+, TESS, and CHEOPS observations highlights the potential of combined studies with space- and ground-based instruments and represents a promising avenue for advancing our understanding of exoplanet atmospheres.
@@ 2 Observations
We observed the thermal emission spectrum of WASP-178b on 25 May 2023 between 04:35 UT and 08:13 UT with CRIRES+ (Dorn et al. 2023). The CRIRES+ instrument is a cryogenic cross-dispersed high-resolution near-infrared spectrograph, installed at the 8-meter Unit Telescope 3 of the Very Large Telescope. The observations of the planet cover the orbital phase interval 0.52–0.57, when the dayside of WASP-178b is almost completely aligned with the observer’s line of sight. A total of 52 consecutive spectra were acquired with an integration time of 240 s. We used an ABBA nodding pattern, which consists of observing the target at two different slit positions (A and B), to facilitate the removal of the sky background and detector artifacts in the subsequent data reduction steps. The slit width was set to 0.2′′, giving a nominal resolving power of 100 000 under the condition of a homogeneous slit illumination. In our observations, however, the performance of the adaptive optics system was very good, resulting in a point spread function (PSF) smaller than the slit, and thus an increased spectral resolution. From the width of the PSF, we determined that the spectral resolution of our observations was 135 000. Our spectra were acquired in the K2166 wavelength setting, which offers a wavelength coverage between 1921 nm and 2472 nm. This setting consists of seven echelle orders, each divided by two narrow gaps between detectors, yielding a total of 21 wavelength segments. The airmass varied between 1.1 and 1.8 from the start to the end of the observations and the mean signal-to-noise (S/N) ratio per detector pixel was 52 (Fig. 1).
@@ 3.1 Extraction of the one-dimensional spectra
To account for the shifts in the spectral data points relative to the wavelength solutions of the pipeline, an additional refinement step was applied. We calculated the mean spectrum at each nodding position and applied molecfit (Smette et al. 2015) to the resulting A and B master spectra. By fitting the telluric absorption lines in the master spectra, molecfit corrected the input wavelength solutions by the respective shift and provided us with an improved solution for the A and B spectra as output. Although the applied procedure results in different molecfit wavelength solutions for A and B, we refrained from shifting the spectral data points to a common wavelength solution to avoid introducing spurious signals during the required interpolation process. Instead, following Yan et al. (2023), we treated the A and B spectra as two independent data sets throughout the subsequent procedures of this section, and combined their information without the need for interpolation in Sects. 4 and 5 after data reduction was complete.
@@ 4 Detection of the planetary emission lines
We used the cross-correlation method (e.g., Snellen et al. 2010; Brogi et al. 2012; Birkby et al. 2013) to extract the faint spectral emission lines of WASP-178b from the noise-dominated residual spectra. This method consists of calculating the cross-correlation function (CCF) between an exoplanet model spectrum and the residual spectra, thereby mapping the information of the faint planetary spectral lines onto a detectable signal peak. We searched for the spectral signature of CO, H2O, and OH, which are the chemical species with the most prominent emission lines expected in the K-band for any reasonable atmospheric composition. In the specific case of CO, the cross-correlation analysis was performed individually on the molecules with the two isotopologues 12CO and 13CO. In addition, our search included the emission signal of Fe i, which has recently been detected with the same CRIRES+ setting as used in this work in the atmosphere of another UHJ, MASCARA-1b (Ramkumar et al. 2023).
@@ 4.1 Model spectra
The - profile is parametrized by a low pressure point (, ) and a high pressure point (, ), with an isothermal atmosphere at pressures below or higher than (Fig. 4). The temperature between the two isothermal layers was assumed to change linearly as a function of . Assuming equilibrium chemistry and the elemental abundances (Asplund et al. 2009) retrieved in Sect. 5, we used FastChem (Stock et al. 2018)333https://github.com/exoclime/FastChem to obtain the mean molecular weight and volume mixing ratio (VMR) of the chemical species studied (Fig. 4). In addition, for each chemical species we needed a --wavelength grid of its opacities to calculate the corresponding model spectrum. The data used to compute the high-resolution opacities were obtained from the following sources: CO line information from Li et al. (2015), H2O data from the POKAZATEL line list (Polyansky et al. 2018), OH opacities from the MoLLIST database (Brooke et al. 2016), Fe i atomic line data from Kurucz (2018). We ran the radiative transfer code petitRADTRANS (Mollière et al. 2019) to generate the high-resolution model spectra of the different chemical species ( ). Finally, the resulting model spectra were convolved with the rotational broadening profile described in Sect. 5.
@@ 5.1.2 Joint retrieval with photometric data
We used petitRADTRANS to calculate the thermal emission component of Eq. 6 over the wavelength range of TESS and CHEOPS. The photometric bandpasses of both instruments cover a significant fraction of the visible wavelength range where strong emission lines of atomic metals, their oxides, and hydrides are expected. In our radiative transfer calculations, we included the opacities of Fe, Ti, Na, K, Ca, Mg, TiO, VO, FeH, CrH, and CaH, which represent the metal species with the most prominent emission features at visible wavelengths for any reasonable atmospheric composition (e.g., Hubeny et al. 2003; Fortney et al. 2008; Lothringer & Barman 2019). Additionally, we accounted for the contribution from the continuum opacity source H-. We provided petitRADTRANS with the same thermal and chemical parametrizations as detailed in Sect. 5.1.1. The resulting emission flux of each instrument was integrated over its respective bandpass using the rebin-give-width tool of petitRADTRANS. We integrated over the range 0.60 to 1.00 m for TESS, while for CHEOPS the integration was performed over the range 0.35 to 1.10 m. Since the model spectrum was integrated over the relatively wide TESS and CHEOPS bandpasses, we used the petitRADTRANS low-resolution mode to speed up the radiative transfer calculations. The opacities used in the low-resolution mode require less dense sampling in wavelength space than those used to generate the high-resolution CRIRES+ model spectra, and are available in the petitRADTRANS opacity database 444https://petitradtrans.readthedocs.io/en/latest/content/available_opacities.html.
@@ 6 Conclusions
We have studied the dayside emission spectrum of the UHJ WASP-178b by analyzing high-resolution spectral observations with CRIRES+ in the K-band. Using the cross-correlation technique and a Bayesian retrieval framework that includes space-based photometric measurements, the following main results were obtained:
@@
We identified the spectral signatures of the two molecular species 12CO and H2O. These are the chemical species with the most dominant emission lines in the spectral K-band, and are among the most important carbon- and oxygen-bearing species. No significant signature of the chemical species 13CO, OH, and Fe i could be identified in our high-resolution spectroscopy data (Sect. 4.3).
@@ Acknowledgements.
CRIRES+ is an ESO upgrade project carried out by Thüringer Landessternwarte Tautenburg, Georg-August Universität Göttingen, and Uppsala University. The project is funded by the Federal Ministry of Education and Research (Germany) through Grants 05A11MG3, 05A14MG4, 05A17MG2 and the Knut and Alice Wallenberg Foundation. This project is based on observations collected at the European Organisation for Astronomical Research in the Southern Hemisphere under the ESO programme 111.254J.001. D.C. is supported by the LMU-Munich Fraunhofer-Schwarzschild Fellowship and by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy - EXC 2094 - 390783311. F.Y. acknowledges the support by the National Natural Science Foundation of China (grant No. 42375118). F.L. acknowledges the support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project number 314665159. S.C. and M.R acknowledge the support of the DFG priority program SPP 1992 “Exploring the Diversity of Extrasolar Planets” (CZ 222/5-1, RE 1664/16-1, and DFG PR 36 24602/41). N.P., L.B.-Ch., and A.D.R. acknowledge support by the Knut and Alice Wallenberg Foundation (grant 2018.0192). E.N. acknowledges the support by the DFG Research Unit FOR2544 “Blue Planets around Red Stars”. D.S. acknowledges financial support from the project PID2021-126365NB-C21(MCI/AEI/FEDER, UE) and from the Severo Ochoa grant CEX2021-001131-S funded by MCIN/AEI/10.13039/501100011033. This work has made use of the following Python packages: Astropy (Astropy Collaboration et al. 2013), CMasher (van der Velden 2020), corner (Foreman-Mackey 2016), GNU Parallel (Tange 2021), Matplotlib (Hunter 2007), NumPy (Harris et al. 2020), PyAstronomy (Czesla et al. 2019), and SciPy (Virtanen et al. 2020). The authors would also like to thank Paul Mollière for his assistance in selecting the molecular line lists and calculating the high-resolution opacity files for use in petitRADTRANS.References

### v3

@@
Despite recent progress in the spectroscopic characterization of individual exoplanets, the atmospheres of key ultra-hot Jupiters (UHJs) still lack comprehensive investigations. These include WASP-178b, one of the most irradiated UHJs known to date. We observed the dayside emission signal of this planet with CRIRES+ in the spectral K band. By applying the cross-correlation technique and a Bayesian retrieval framework to the high-resolution spectra, we identified the emission signature of 12CO (S/N = 8.9) and H2O (S/N = 4.9), and a strong atmospheric thermal inversion. A joint retrieval with space-based secondary eclipse measurements from TESS and CHEOPS allowed us to refine our results on the thermal profile and thus to constrain the atmospheric chemistry, yielding a solar to super-solar metallicity ( dex) and a solar C/O ratio (). We infer a significant excess of spectral line broadening and identify a slight Doppler-shift between the 12CO and H2O signals. These findings provide strong evidence for a super-rotating atmospheric flow pattern and suggest the possible existence of chemical inhomogeneities across the planetary dayside hemisphere. In addition, the inclusion of photometric data in our retrieval allows us to account for stellar light reflected by the planetary atmosphere, resulting in an upper limit on the geometric albedo (0.23). The successful characterization of WASP-178b’s atmosphere through a joint analysis of CRIRES+, TESS, and CHEOPS observations highlights the potential of combined studies with space- and ground-based instruments and represents a promising avenue for advancing our understanding of exoplanet atmospheres.
@@ 2 Observations
We observed the thermal emission spectrum of WASP-178b on 25 May 2023 between 04:35 UT and 08:13 UT with CRIRES+ (Dorn et al. 2023). The CRIRES+ instrument is a cryogenic cross-dispersed high-resolution near-infrared spectrograph, installed at the 8-meter Unit Telescope 3 of the Very Large Telescope. The observations of the planet cover the orbital phase interval 0.52–0.57, when the dayside of WASP-178b is almost completely aligned with the observer’s line of sight. A total of 52 consecutive spectra were acquired with an integration time of 240 s. We used an ABBA nodding pattern, which consists of observing the target at two different slit positions (A and B), to facilitate the removal of the sky background and detector artifacts in the subsequent data reduction steps. The slit width was set to 0.2′′, giving a nominal resolving power of 100 000 under the condition of a homogeneous slit illumination. In our observations, however, the performance of the adaptive optics system was very good, resulting in a point spread function (PSF) smaller than the slit, and thus an increased spectral resolution. From the width of the PSF, we determined that the spectral resolution of our observations was 135 000. Our spectra were acquired in the K2166 wavelength setting, which offers a wavelength coverage between 1921 nm and 2472 nm. This setting consists of seven echelle orders, each divided by two narrow gaps between detectors, yielding a total of 21 wavelength segments. The airmass varied between 1.1 and 1.8 from the start to the end of the observations and the mean signal-to-noise ratio (S/N) per detector pixel was 52 (Fig. 1).
@@ 3.1 Extraction of the one-dimensional spectra
To account for the shifts in the spectral data points relative to the wavelength solutions of the pipeline, an additional refinement step was applied. We calculated the mean spectrum at each nodding position and applied molecfit (Smette et al. 2015) to the resulting A and B master spectra. By fitting the telluric absorption lines in the master spectra, molecfit corrected the input wavelength solutions by the respective shift and provided us with an improved solution for the A and B spectra as output333The extracted data, together with the refined wavelength solution, are available at https://zenodo.org/records/11637332 (Lavail 2024). Although the applied procedure results in different molecfit wavelength solutions for A and B, we refrained from shifting the spectral data points to a common wavelength solution to avoid introducing spurious signals during the required interpolation process. Instead, following Yan et al. (2023), we treated the A and B spectra as two independent data sets throughout the subsequent procedures of this section, and combined their information without the need for interpolation in Sects. 4 and 5 after data reduction was complete.
@@ 4 Detection of the planetary emission lines
We used the cross-correlation method (e.g., Snellen et al. 2010; Brogi et al. 2012; Birkby et al. 2013) to extract the faint spectral emission lines of WASP-178b from the noise-dominated residual spectra. This method consists of calculating the cross-correlation function (CCF) between an exoplanet model spectrum and the residual spectra, thereby mapping the information of the faint planetary spectral lines onto a detectable signal peak. We searched for the spectral signature of CO, H2O, and OH, which are the chemical species with the most prominent emission lines expected in the K band for any reasonable atmospheric composition. In the specific case of CO, the cross-correlation analysis was performed individually on the molecules with the two isotopologues 12CO and 13CO. In addition, our search included the emission signal of Fe i, which has recently been detected with the same CRIRES+ setting as used in this work in the atmosphere of another UHJ, MASCARA-1b (Ramkumar et al. 2023).
@@ 4.1 Model spectra
The - profile is parametrized by a low pressure point (, ) and a high pressure point (, ), with an isothermal atmosphere at pressures below or higher than (Fig. 4). The temperature between the two isothermal layers was assumed to change linearly as a function of . Assuming equilibrium chemistry and the elemental abundances (Asplund et al. 2009) retrieved in Sect. 5, we used FastChem (Stock et al. 2018)444https://github.com/exoclime/FastChem to obtain the mean molecular weight and volume mixing ratio (VMR) of the chemical species studied (Fig. 4). In addition, for each chemical species we needed a --wavelength grid of its opacities to calculate the corresponding model spectrum. The data used to compute the high-resolution opacities were obtained from the following sources: CO line information from Li et al. (2015), H2O data from the POKAZATEL line list (Polyansky et al. 2018), OH opacities from the MoLLIST database (Brooke et al. 2016), Fe i atomic line data from Kurucz (2018). We ran the radiative transfer code petitRADTRANS (Mollière et al. 2019) to generate the high-resolution model spectra of the different chemical species ( ). Finally, the resulting model spectra were convolved with the rotational broadening profile described in Sect. 5.
@@ 5.1.2 Joint retrieval with photometric data
We used petitRADTRANS to calculate the thermal emission component of Eq. 6 over the wavelength range of TESS and CHEOPS. The photometric bandpasses of both instruments cover a significant fraction of the visible wavelength range where strong emission lines of atomic metals, their oxides, and hydrides are expected. In our radiative transfer calculations, we included the opacities of Fe, Ti, Na, K, Ca, Mg, TiO, VO, FeH, CrH, and CaH, which represent the metal species with the most prominent emission features at visible wavelengths for any reasonable atmospheric composition (e.g., Hubeny et al. 2003; Fortney et al. 2008; Lothringer & Barman 2019). Additionally, we accounted for the contribution from the continuum opacity source H-. We provided petitRADTRANS with the same thermal and chemical parametrizations as detailed in Sect. 5.1.1. The resulting emission flux of each instrument was integrated over its respective bandpass using the rebin-give-width tool of petitRADTRANS. We integrated over the range 0.60 to 1.00 m for TESS, while for CHEOPS the integration was performed over the range 0.35 to 1.10 m. Since the model spectrum was integrated over the relatively wide TESS and CHEOPS bandpasses, we used the petitRADTRANS low-resolution mode to speed up the radiative transfer calculations. The opacities used in the low-resolution mode require less dense sampling in wavelength space than those used to generate the high-resolution CRIRES+ model spectra, and are available in the petitRADTRANS opacity database 555https://petitradtrans.readthedocs.io/en/latest/content/available_opacities.html.
@@ 6 Conclusions
We have studied the dayside emission spectrum of the UHJ WASP-178b by analyzing high-resolution spectral observations with CRIRES+ in the K band. Using the cross-correlation technique and a Bayesian retrieval framework that includes space-based photometric measurements, the following main results were obtained:
@@
We identified the spectral signatures of the two molecular species 12CO and H2O. These are the chemical species with the most dominant emission lines in the spectral K band, and are among the most important carbon- and oxygen-bearing species. No significant signature of the chemical species 13CO, OH, and Fe i could be identified in our high-resolution spectroscopy data (Sect. 4.3).
@@ Acknowledgements.
The authors thank the referee for very useful comments and suggestions. CRIRES+ is an ESO upgrade project carried out by Thüringer Landessternwarte Tautenburg, Georg-August Universität Göttingen, and Uppsala University. The project is funded by the Federal Ministry of Education and Research (Germany) through Grants 05A11MG3, 05A14MG4, 05A17MG2 and the Knut and Alice Wallenberg Foundation. This project is based on observations collected at the European Organisation for Astronomical Research in the Southern Hemisphere under the ESO programme 111.254J.001. D.C. is supported by the LMU-Munich Fraunhofer-Schwarzschild Fellowship and by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy - EXC 2094 - 390783311. F.Y. acknowledges the support by the National Natural Science Foundation of China (grant No. 42375118). F.L. acknowledges the support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project number 314665159. S.C. and M.R acknowledge the support of the DFG priority program SPP 1992 “Exploring the Diversity of Extrasolar Planets” (CZ 222/5-1, RE 1664/16-1, and DFG PR 36 24602/41). N.P., L.B.-Ch., and A.D.R. acknowledge support by the Knut and Alice Wallenberg Foundation (grant 2018.0192). E.N. acknowledges the support by the DFG Research Unit FOR2544 “Blue Planets around Red Stars”. D.S. acknowledges financial support from the project PID2021-126365NB-C21(MCI/AEI/FEDER, UE) and from the Severo Ochoa grant CEX2021-001131-S funded by MCIN/AEI/10.13039/501100011033. This work has made use of the following Python packages: Astropy (Astropy Collaboration et al. 2013), CMasher (van der Velden 2020), corner (Foreman-Mackey 2016), GNU Parallel (Tange 2021), Matplotlib (Hunter 2007), NumPy (Harris et al. 2020), PyAstronomy (Czesla et al. 2019), and SciPy (Virtanen et al. 2020). The authors would also like to thank Paul Mollière for his assistance in selecting the molecular line lists and calculating the high-resolution opacity files for use in petitRADTRANS.References
@@
- Lavail (2024) Lavail, A. 2024, CRIRES+ reduced spectroscopic observations of WASP-178